# AI Software Developers as Generalist Agents

### v2

@@ Abstract
[Uncaptioned image] Benchmark https://hf.co/spaces/OpenHands/evaluation
@@ 2.3 Agent Skills: The Extensible Agent-Computer Interface
Rigorously tested and maintained. We follow best practices in software engineering and write extensive unit tests for tools in AgentSkills to ensure their reliability and usability.
@@ 4.2 Software Engineering
SWE-Bench (Jimenez et al., 2024) is designed to assess agents’ abilities in solving real-world GitHub issues, such as bug reports or feature requests. The agent interacts with the repository and attempts to fix the issue provided through file editing and code execution. The agent-modified code repository is tested against a test suite incorporating new tests added from human developers’ fixes for the same issue. Each test instance accompanies a piece of “hint text” that consists of natural language suggestions for how to solve the problem. Throughout this paper, we report all results without using hint text. A canonical subset, SWE-bench Lite, is created to facilitate accessible and efficient testing. We default to use this subset for testing for cost-saving consideration. [^2] Running the complete set of 2294 instances costs \$6.9k, using a conservative estimate of \$3 per instance. Result. As shown in Tab. 4, our most recent version of CodeActAgent v1.8, using claude-3.5-sonnet, achieves a competitive resolve rate of $26\%$ compared to other open-source software development specialists.
@@
Agent Model Success Rate (%) \$ Avg. Cost SWE-Bench Lite (Jimenez et al., 2024), 300 instances, w/o Hint SWE-Agent (Yang et al., 2024) gpt-4-1106-preview $18.0$ $1.67$ AutoCodeRover (Zhang et al., 2024b) gpt-4-0125-preview $19.0$ - Aider (Gauthier, ) gpt-4o & claude-3-opus $26.3$ - OH CodeActAgent v1.8 gpt-4o-mini-2024-07-18 $7.0$ $0.01$ gpt-4o-2024-05-13 $22.0$ $1.72$ claude-3-5-sonnet@20240620 $26.0$ $1.10$ HumanEvalFix (Muennighoff et al., 2024), 164 instances Prompting, 0-shot BLOOMZ-176B $16.6$ - OctoCoder-15B $30.4$ - DeepSeekCoder-33B-Instruct $47.5$ - StarCoder2-15B $48.6$ - SWE-agent, 1-shot (Yang et al., 2024) gpt-4-turbo $87.7$ - OH CodeActAgent v1.5, Generalist, 0-shot. gpt-3.5-turbo-16k-0613 $20.1$ $0.11$ gpt-4o-2024-05-13 $79.3$ $0.14$ BIRD (Li et al., 2023b), 300 instances Prompting, 0-shot CodeLlama-7B-Instruct $18.3$ - CodeQwen-7B-Chat $31.3$ - OH CodeActAgent v1.5 gpt-4-1106-preview $42.7$ $0.19$ gpt-4o-2024-05-13 $47.3$ $0.11$ ML-Bench (Tang et al., 2024b), 68 instances prompting + BM25, 0-shot gpt-3.5-turbo $11.0$ - gpt-4-1106-preview $22.1$ - gpt-4o-2024-05-13 $26.2$ - SWE-Agent (Yang et al., 2024) gpt-4-1106-preview $42.6$ $1.91$ Aider (Gauthier, ) gpt-4o $64.4$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $76.5$ $0.25$ gpt-4-1106-preview $58.8$ $1.22$ gpt-3.5-turbo-16k-0613 $13.2$ $0.12$ BioCoder (Python) (Tang et al., 2024b), 157 instances prompting, 0-shot gpt-3.5-turbo $11.0$ - gpt-4-1106-preview $12.7$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $27.5$ $0.13$ BioCoder (Java) (Tang et al., 2024b), 50 instances prompting, 0-shot gpt-3.5-turbo $4.1$ - gpt-4-1106-preview $6.4$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $44.0$ $0.11$ Gorilla APIBench (Patil et al., 2023), 1775 instances Prompting, 0-shot claude-v1 $8.7$ - gpt-4-0314 $21.2$ - gpt-3.5-turbo-0301 $29.7$ - Gorilla, finetuned for API calls, 0-shot (Patil et al., 2023; Touvron et al., 2023) llama-7b 75.0 - OH CodeActAgent v1.5 gpt-3.5-turbo-0125 $21.6$ $0.002$ gpt-4o-2024-05-13 $36.4$ $0.04$ ToolQA (Zhuang et al., 2024), 800 instances Prompting, 0-shot ChatGPT + CoT $5.1$ - ChatGPT $5.6$ - Chameleon $10.6$ - ReAct, 0-shot (Yao et al., 2023; OpenAI, 2024a) gpt-3.5-turbo $36.8$ - gpt-3 $43.1$ - OH CodeActAgent v1.5 gpt-3.5-turbo-0125 $2.3$ $0.03$ gpt-4o-2024-05-13 $47.2$ $0.91$
@@ 4.2.1 HumanEvalFix
ML-Bench (Tang et al., 2024b) evaluates agents’ ability to solve machine learning tasks across 18 GitHub repositories. The benchmark comprises 9,641 tasks spanning 169 diverse ML problems, requiring agents to generate bash scripts or Python code in response to user instructions. In the sandbox environment, agents can iteratively execute commands and receive feedback, allowing them to understand the repository context and fulfill user requirements progressively. Following the setup from the original paper, we perform agent evaluation on the quarter subset of ML-Bench. Results. As shown in Table 4, OpenHands agents with GPT-4o achieve the highest success rate of $76.47\%$ on ML-Bench, outperforming SWE-Agent ($42.64\%$). Performance drops with less capable models. These results demonstrate the effectiveness of OpenHands agent in complex ML tasks. We notice that agents show potential in reducing hallucination and syntax errors compared to non-agent approaches in the ML-LLM-Bench settings (Tang et al., 2024b).
@@
Gorilla APIBench (Patil et al., 2023) evaluates agents’ abilities to use APIs. it incorporates tasks on TorchHub, TensorHub, and HuggingFace. During the evaluation, models are given a question related to API usage, such as "identify an API capable of converting spoken language in a recording to text." Correctness is evaluated based on whether the model’s API call is in the correct domain. Results. As shown in Table 4, OpenHands using GPT-4o, with a success rate of $36.4\%$, outperforms baselines not specifically finetuned for API calling. While Gorilla shows higher performance on APIBench, Patil et al. (2023) finetune this model for API calling in particular.
@@
ToolQA (Zhuang et al., 2024) evaluates agents’ abilities to use external tools. This benchmark includes tasks on various topics like flight status, coffee price, Yelp data, and Airbnb data, requiring the use of various tools such as text tools, database tools, math tools, graph tools, code tools, and system tools. It features two levels: easy and hard. Easy questions focus more on single-tool usage, while hard questions emphasize reasoning. We adopt the easy subset for evaluation. Results. Compared to all baselines, OpenHands with GPT-4o shows the highest performance. We notice that agents perform better on tasks related to CSV and database tool usage but requires improvements on math and calculator tool usage.
@@
BioCoder (Tang et al., 2024c) is a repository-level code generation benchmark that evaluates agents’ performance on bioinformatics-related tasks, specifically the ability to retrieve and accurately utilize context. The original prompts contain the relevant context of the code; however, in this study, we have removed them to demonstrate the capability of OpenHands to perform context retrieval, self-debugging, and reasoning in multi-turn interactions. BioCoder consists of 157 Python and 50 Java functions, each targeting a specific area in bioinformatics, such as proteomics, genomics, and other specialized domains. The benchmark targets real-world code by generating code in existing repositories where the relevant code has been masked out. Results. Table 4 shows that OpenHands, using GPT-4o, achieves a success rate of $44.0\%$. This outperforms all prompting-based non-agent baselines, with GPT-4 alone only achieving $6.4\%$.
@@
BIRD (Li et al., 2023b) is a benchmark for text-to-SQL tasks (i.e., translate natural language into executable SQL) aimed at realistic and large-scale database environments. We select 300 samples from the dev set to integrate into OpenHands and evaluate on execution accuracy. Additionally, we extend the setting by allowing the agent to engage in multi-turn interactions to arrive at the final SQL query, enabling it to correct historical results by observing the results of SQL execution. Results. As shown in Table 4, OpenHands with GPT-4o achieves an execution accuracy of $47.3\%$ on a subset of BIRD, showcasing the potential of OpenHands as a SQL agent. The result outperforms approaches utilizing prompting with code LLMs, such as CodeLlama-7B-Instruct ($18.3\%$) and CodeQwen-7B-Chat (Bai et al., 2023) ($31.3\%$).
@@ 4.3 Web Browsing
WebArena (Zhou et al., 2023a) is a self-hostable, execution-based web agent benchmark that allows agents to freely choose which path to take in completing their given tasks. WebArena comprises 812 human-curated task instructions across various domains, including shopping, forums, developer platforms, and content management systems. Each task is paired with a handwritten test case that verifies agent success, e.g., by checking the status of a web page element against a reference or the textual answer returned by the agent. Results. From Tab. 5, we can see that our BrowsingAgent achieves competitive performance among agents that use LLMs with domain-general prompting techniques. Some agents (e.g., AutoWebGLM) require manual effort tailored to the WebArena task domain. This showcases the performance trade-off between a generalist vs. a domain-tailored specialist web agent, and we opt for a more general browsing agent as a building block in OpenHands.
MiniWoB++ (Liu et al., 2018) is an interactive web benchmark, with built-in reward functions. The tasks are synthetically initialized on 125 different minimalist web interfaces. Unlike WebArena, tasks are easier without page changes, require fewer steps, and provide low-level step-by-step task directions. Note that it contains a portion of environments that require vision capability to tackle successfully, and many existing work choose to focus only on a subset of the tasks (Kim et al., 2024; Li et al., 2023d; Shaw et al., 2023). Still, we report the performance on the full set and only include baselines that are evaluated on the full set. Results. From Tab. 5, we see that our BrowsingAgent finishes nearly half of the tasks without any adaptation to the environment. However, due to the synthetic nature of MiniWoB++, the state-of-the-art agents explicitly trained for the environments with reinforcement learning and/or human behavior cloning have almost saturated the performance.
@@
Agent Model Success Rate (%) \$ Avg. Cost WebArena (Zhou et al., 2023a), 812 instances Lemur (Xu et al., 2023) Lemur-chat-70b $5.3$ - Patel et al. (2024) Trained 72B with self-improvement synthetic data $9.4$ - AutoWebGLM (Lai et al., 2024) Trained 7B with human/agent hybrid annotation $18.2$ - Auto Eval & Refine (Pan et al., 2024) GPT-4 + Reflexion w/ GPT-4V reward model $20.2$ - WebArena Agent (Zhou et al., 2023a) Llama3-chat-8b $3.3$ - Llama3-chat-70b $7.0$ - gpt-3.5-turbo $6.2$ - gpt-4-turbo $14.4$ - OH BrowsingAgent v1.0 gpt-3.5-turbo-0125 $5.2$ $0.02$ gpt-4o-mini-2024-07-18 $8.5$ $0.01$ gpt-4o-2024-05-13 $14.8$ $0.15$ claude-3-5-sonnet-20240620 $15.5$ $0.10$ OH CodeActAgent v1.8 via delegation to BrowsingAgent v1.0 gpt-4o-mini-2024-07-18 $8.3$ - gpt-4o-2024-05-13 $14.5$ - claude-3-5-sonnet-20240620 $15.3$ - MiniWoB++ (Liu et al., 2018), 125 environments Workflow Guided Exploration (Liu et al., 2018) Trained specialist model with environment exploration $34.6$ - CC-NET (Humphreys et al., 2022) Trained specialist model with RL and human annotated BC $91.1$ - OH BrowsingAgent v1.0 gpt-3.5-turbo-0125 $27.2$ $0.01$ gpt-4o-2024-05-13 $40.8$ $0.05$ OH CodeActAgent v1.8 via delegation to BrowsingAgent v1.0 gpt-4o-2024-05-13 $39.8$ -
@@ 4.4 Miscellaneous Assistance
GAIA (Mialon et al., 2023) evaluates agents’ general task-solving skills, covering different real-world scenarios. It requires various agent capabilities, including reasoning, multi-modal understanding, web browsing, and coding. GAIA consists of 466 curated tasks across three levels. Setting up GAIA is traditionally challenging due to the complexity of integrating various tools with the agent, but OpenHands’s infrastructure (e.g., runtime §2.2, tools §2.3) simplifies the integration significantly. Results. In our experiments, we achieved a score of $32.1$ on the GAIA (level-1 val), significantly improving over the original AutoGPT (Gravitas, 2023). GAIA is sensitive to the support of multimodal input and web navigation skills, suggesting further score improvements as OpenHands’s infrastructure improves.
GPQA (Rein et al., 2023) evaluates agents’ ability for coordinated tool use when solving challenging graduate-level problems. It consists of 448 curated and difficult multiple-choice questions in biology, physics, and chemistry. Tool use (e.g., python) and web search are often useful to assist agents in answering these questions since they provide accurate calculations that LLMs are often incapable of and access to information outside of the LLM’s parametric knowledge base. Results. Results are shown in Tab. 6 and 7. We observe that OpenHands’s integrated for supporting diverse tool use (e.g., python for calculations) as well as web-search (for searching relevant facts) allows the resulting agent to better solve complex multi-step problems, surpassing the prior state-of-the-art by 9.6% and 12.3% on the main and diamond subsets respectively on GPQA (Rein et al., 2023).
@@
Agent Model Success Rate (%) \$ Avg. Cost GAIA (Mialon et al., 2023), L1 validation set, 53 instances AutoGPT (Gravitas, 2023) gpt-4-turbo $13.2$ - OH GPTSwarm v1.0 gpt-4-0125-preview $30.2$ $0.110$ gpt-4o-2024-05-13 $32.1$ $0.050$ GPQA (Rein et al., 2023), diamond set, 198 instances (refer to §G, Tab. 7 for other subsets) Human (Rein et al., 2023) Expert human $81.3$ - Non-expert human $21.9$ - Few-shot Prompting + Chain-of-Thought (Rein et al., 2023) gpt-3.5-turbo-16k $29.6$ - gpt-4 $38.8$ - OH CodeActAgent v1.8 claude-3-5-sonnet-20240620 $52.0$ $0.065$ AgentBench (Liu et al., 2023), OS (bash) subset, 144 instances AgentBench Baseline Agent (Liu et al., 2023) gpt-4 $42.4$ - gpt-3.5-turbo $32.6$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $57.6$ $0.085$ gpt-3.5-turbo-0125 $11.8$ $0.006$ MINT (Wang et al., 2024b): math subset, 225 instances MINT Baseline Agent gpt-4-0613 $65.8$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $77.3$ $0.070$ gpt-3.5-turbo-16k-0613 $33.8$ $0.048$ MINT (Wang et al., 2024b): code subset, 136 instances MINT Baseline Agent gpt-4-0613 $59.6$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $50.0$ $0.087$ gpt-3.5-turbo-16k-0613 $5.2$ $0.030$ ProofWriter (Tafjord et al., 2021), 600 instances Few-shot Prompting + Chain-of-Thought (Pan et al., 2023) gpt4 $68.1$ - Logic-LM (Pan et al., 2023) gpt4 + symbolic solver $79.6$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $78.8$ - Entity Deduction Arena (Zhang et al., 2024a), 200 instances Human - $21.0$ - Zero-shot Prompting (Zhang et al., 2024a) gpt-4-0314 $40.0$ - gpt-3.5-turbo-0613 $27.0$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $38.0$ - gpt-3.5-turbo-16k-0613 $24.0$ -
@@
AgentBench (Liu et al., 2023) evaluates agents’ reasoning and decision-making abilities in a multi-turn, open-ended generation setting. We selected the code-grounded operating system (OS) subset with 144 tasks. Agents from OpenHands interact directly with the task-specific OS using bash commands in a multi-turn manner, combining interaction and reasoning to automate task completion. Results. In our experiments (Tab. 6), OpenHands CodeActAgent v1.5 achieves a score of $57.6\%$ on the AgentBench using gpt-4o, outperforming the $42.4\%$ baseline using gpt-4 from the original paper. Interestingly, when employing weaker models such as gpt-3.5-turbo, OpenHands agents generally underperform compared to the original baseline agents. This finding suggests that generalist agents, like those implemented in OpenHands, require a certain threshold of foundation model capability - particularly instruction following - to function effectively.
@@
MINT (Wang et al., 2024b) is a benchmark designed to evaluate agents’ ability to solve challenging tasks through multi-turn interactions using tools and natural language feedback simulated by GPT-4. We use coding and math subsets used in Yuan et al. (2024). We follow the original paper and allow the agent to interact with up to five iterations with two chances to propose solutions. Results. As shown in Tab. 6), OpenHands agents achieve comparable performance to the default agent in the original benchmark, with a performance improvement in the math subset.
@@
ProofWriter (Tafjord et al., 2021) is a synthetic dataset created to assess deductive reasoning abilities of LLMs. Same as Logic-LM (Pan et al., 2023), we focus on the most challenging subset, which contains 600 instances requiring 5-hop reasoning. To minimize the impact of potential errors in semantic parsing, we use the logical forms provided by Logic-LM. Results. In Tab. 6, OpenHands agent employs a symbolic solver to solve the task, achieving performance comparable to the state-of-the-art neuro-symbolic model (i.e., Logic-LM) (Pan et al., 2023).
@@
Entity Deduction Arena (EDA) (Zhang et al., 2024a) evaluates agents’ ability to deduce unknown entities through strategic questioning, akin to the 20 Questions game. This benchmark tests the agent’s state tracking, strategic planning, and inductive reasoning capabilities over multi-turn conversations. We evaluate two datasets “Things” and “Celebrities”, each comprising 100 instances, and report the average success rate over these two datasets. Results. Tab. 6 shows that CodeActAgent yields comparable performance comparing with the results reported in the original paper (Zhang et al., 2024a).

### v3

@@ 4.2 Software Engineering
SWE-Bench (Jimenez et al., 2024) is designed to assess agents’ abilities in solving real-world GitHub issues, such as bug reports or feature requests. The agent interacts with the repository and attempts to fix the issue provided through file editing and code execution. The agent-modified code repository is tested against a test suite incorporating new tests added from human developers’ fixes for the same issue. Each test instance accompanies a piece of “hint text” that consists of natural language suggestions for how to solve the problem. Throughout this paper, we report all results without using hint text. A canonical subset, SWE-bench Lite, is created to facilitate accessible and efficient testing. We default to use this subset for testing for cost-saving consideration. [^2] Running the complete set of 2294 instances costs \$6.9k, using a conservative estimate of \$3 per instance. Result. As shown in Tab. 4, our most recent version of CodeActAgent v1.8, using claude-3.5-sonnet, achieves a competitive resolve rate of $26\%$ compared to other open-source SWE specialists.
@@
Agent Model Success Rate (%) \$ Avg. Cost SWE-Bench Lite (Jimenez et al., 2024), 300 instances, w/o Hint SWE-Agent (Yang et al., 2024) gpt-4-1106-preview $18.0$ $1.67$ AutoCodeRover (Zhang et al., 2024b) gpt-4-0125-preview $19.0$ - Aider (Gauthier, ) gpt-4o & claude-3-opus $26.3$ - OH CodeActAgent v1.8 gpt-4o-mini-2024-07-18 $7.0$ $0.01$ gpt-4o-2024-05-13 $22.0$ $1.72$ claude-3-5-sonnet@20240620 $26.0$ $1.10$ HumanEvalFix (Muennighoff et al., 2024), 164 instances Prompting, 0-shot BLOOMZ-176B $16.6$ - OctoCoder-15B $30.4$ - DeepSeekCoder-33B-Instruct $47.5$ - StarCoder2-15B $48.6$ - SWE-agent, 1-shot (Yang et al., 2024) gpt-4-turbo $87.7$ - OH CodeActAgent v1.5, Generalist, 0-shot. gpt-3.5-turbo-16k-0613 $20.1$ $0.11$ gpt-4o-2024-05-13 $79.3$ $0.14$ BIRD (Li et al., 2023b), 300 instances Prompting, 0-shot CodeLlama-7B-Instruct $18.3$ - CodeQwen-7B-Chat $31.3$ - OH CodeActAgent v1.5 gpt-4-1106-preview $42.7$ $0.19$ gpt-4o-2024-05-13 $47.3$ $0.11$ ML-Bench (Tang et al., 2024b), 68 instances prompting + BM25, 0-shot gpt-3.5-turbo $11.0$ - gpt-4-1106-preview $22.1$ - gpt-4o-2024-05-13 $26.2$ - SWE-Agent (Yang et al., 2024) gpt-4-1106-preview $42.6$ $1.91$ Aider (Gauthier, ) gpt-4o $64.4$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $76.5$ $0.25$ gpt-4-1106-preview $58.8$ $1.22$ gpt-3.5-turbo-16k-0613 $13.2$ $0.12$ BioCoder (Python) (Tang et al., 2024b), 157 instances prompting, 0-shot gpt-3.5-turbo $11.0$ - gpt-4-1106-preview $12.7$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $27.5$ $0.13$ Gorilla APIBench (Patil et al., 2023), 1775 instances Prompting, 0-shot claude-v1 $8.7$ - gpt-4-0314 $21.2$ - gpt-3.5-turbo-0301 $29.7$ - Gorilla, finetuned for API calls, 0-shot (Patil et al., 2023; Touvron et al., 2023) llama-7b 75.0 - OH CodeActAgent v1.5 gpt-3.5-turbo-0125 $21.6$ $0.002$ gpt-4o-2024-05-13 $36.4$ $0.04$ ToolQA (Zhuang et al., 2024), 800 instances Prompting, 0-shot ChatGPT + CoT $5.1$ - ChatGPT $5.6$ - Chameleon $10.6$ - ReAct, 0-shot (Yao et al., 2023; OpenAI, 2024a) gpt-3.5-turbo $36.8$ - gpt-3 $43.1$ - OH CodeActAgent v1.5 gpt-3.5-turbo-0125 $2.3$ $0.03$ gpt-4o-2024-05-13 $47.2$ $0.91$
@@ 4.2.1 HumanEvalFix
ML-Bench (Tang et al., 2024b) evaluates agents’ ability to solve machine learning tasks across 18 GitHub repositories. The benchmark comprises 9,641 tasks spanning 169 diverse ML problems, requiring agents to generate bash scripts or Python code in response to user instructions. In the sandbox environment, agents can iteratively execute commands and receive feedback, allowing them to understand the repository context and fulfill user requirements progressively. Following the setup from the original paper, we perform agent evaluation on the quarter subset of ML-Bench.
@@
Gorilla APIBench (Patil et al., 2023) evaluates agents’ abilities to use APIs. it incorporates tasks on TorchHub, TensorHub, and HuggingFace. During the evaluation, models are given a question related to API usage, such as "identify an API capable of converting spoken language in a recording to text." Correctness is evaluated based on whether the model’s API call is in the correct domain.
@@
ToolQA (Zhuang et al., 2024) evaluates agents’ abilities to use external tools. This benchmark includes tasks on various topics like flight status, coffee price, Yelp data, and Airbnb data, requiring the use of various tools such as text tools, database tools, math tools, graph tools, code tools, and system tools. It features two levels: easy and hard. Easy questions focus more on single-tool usage, while hard questions emphasize reasoning. We adopt the easy subset for evaluation.
@@
BioCoder (Tang et al., 2024c) is a repository-level code generation benchmark that evaluates agents’ performance on bioinformatics-related tasks, specifically the ability to retrieve and accurately utilize context. The original prompts contain the relevant context of the code; however, in this study, we have removed them to demonstrate the capability of OpenHands to perform context retrieval, self-debugging, and reasoning in multi-turn interactions. BioCoder consists of 157 Python and 50 Java functions, each targeting a specific area in bioinformatics, such as proteomics, genomics, and other specialized domains. The benchmark targets real-world code by generating code in existing repositories where the relevant code has been masked out.
@@
BIRD (Li et al., 2023b) is a benchmark for text-to-SQL tasks (i.e., translate natural language into executable SQL) aimed at realistic and large-scale database environments. We select 300 samples from the dev set to integrate into OpenHands and evaluate on execution accuracy. Additionally, we extend the setting by allowing the agent to engage in multi-turn interactions to arrive at the final SQL query, enabling it to correct historical results by observing the results of SQL execution.
@@ 4.3 Web Browsing
WebArena (Zhou et al., 2023a) is a self-hostable, execution-based web agent benchmark that allows agents to freely choose which path to take in completing their given tasks. WebArena comprises 812 human-curated task instructions across various domains, including shopping, forums, developer platforms, and content management systems. Results. From Tab. 5, we can see that our BrowsingAgent achieves competitive performance among agents that use LLMs with domain-general prompting techniques.
MiniWoB++ (Liu et al., 2018) is an interactive web benchmark, with built-in reward functions. The tasks are synthetically initialized on 125 different minimalist web interfaces. Unlike WebArena, tasks are easier without page changes, require fewer steps, and provide low-level step-by-step task directions. Note that it contains a portion of environments that require vision capability to tackle successfully, and many existing work choose to focus only on a subset of the tasks (Kim et al., 2024; Li et al., 2023d; Shaw et al., 2023). Still, we report the performance on the full set and only include baselines that are evaluated on the full set.
@@
Agent Model Success Rate (%) \$ Avg. Cost WebArena (Zhou et al., 2023a), 812 instances Lemur (Xu et al., 2023) Lemur-chat-70b $5.3$ - Patel et al. (2024) Trained 72B with self-improvement synthetic data $9.4$ - AutoWebGLM (Lai et al., 2024) Trained 7B with human/agent hybrid annotation $18.2$ - Auto Eval & Refine (Pan et al., 2024) GPT-4 + Reflexion w/ GPT-4V reward model $20.2$ - WebArena Agent (Zhou et al., 2023a) gpt-3.5-turbo $6.2$ - gpt-4-turbo $14.4$ - OH BrowsingAgent v1.0 gpt-4o-mini-2024-07-18 $8.5$ $0.01$ gpt-4o-2024-05-13 $14.8$ $0.15$ claude-3-5-sonnet-20240620 $15.5$ $0.10$ OH CodeActAgent v1.8 via delegation to BrowsingAgent v1.0 gpt-4o-mini-2024-07-18 $8.3$ - gpt-4o-2024-05-13 $14.5$ - claude-3-5-sonnet-20240620 $15.3$ - MiniWoB++ (Liu et al., 2018), 125 environments Workflow Guided Exploration (Liu et al., 2018) Trained specialist model with environment exploration $34.6$ - CC-NET (Humphreys et al., 2022) Trained specialist model with RL and human annotated BC $91.1$ - OH BrowsingAgent v1.0 gpt-3.5-turbo-0125 $27.2$ $0.01$ gpt-4o-2024-05-13 $40.8$ $0.05$ OH CodeActAgent v1.8 via delegation to BrowsingAgent v1.0 gpt-4o-2024-05-13 $39.8$ -
@@ 4.4 Miscellaneous Assistance
GAIA (Mialon et al., 2023) evaluates agents’ general task-solving skills, covering different real-world scenarios. It requires various agent capabilities, including reasoning, multi-modal understanding, web browsing, and coding. GAIA consists of 466 curated tasks across three levels. Setting up GAIA is traditionally challenging due to the complexity of integrating various tools with the agent, but OpenHands’s infrastructure (e.g., runtime §2.2, tools §2.3) simplifies the integration significantly.
GPQA (Rein et al., 2023) evaluates agents’ ability for coordinated tool use when solving challenging graduate-level problems. Tool use (e.g., python) and web search are often useful to assist agents in answering these questions since they provide accurate calculations that LLMs are often incapable of and access to information outside of the LLM’s parametric knowledge base.
@@
Agent Model Success Rate (%) \$ Avg. Cost GAIA (Mialon et al., 2023), L1 validation set, 53 instances AutoGPT (Gravitas, 2023) gpt-4-turbo $13.2$ - OH GPTSwarm v1.0 gpt-4-0125-preview $30.2$ $0.110$ gpt-4o-2024-05-13 $32.1$ $0.050$ GPQA (Rein et al., 2023), diamond set, 198 instances (refer to §G, Tab. 7 for other subsets) Human (Rein et al., 2023) Expert human $81.3$ - Non-expert human $21.9$ - Few-shot Prompting + Chain-of-Thought (Rein et al., 2023) gpt-3.5-turbo-16k $29.6$ - gpt-4 $38.8$ - OH CodeActAgent v1.8 claude-3-5-sonnet-20240620 $52.0$ $0.065$ AgentBench (Liu et al., 2023), OS (bash) subset, 144 instances AgentBench Baseline Agent (Liu et al., 2023) gpt-4 $42.4$ - gpt-3.5-turbo $32.6$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $57.6$ $0.085$ gpt-3.5-turbo-0125 $11.8$ $0.006$ MINT (Wang et al., 2024b): math subset, 225 instances MINT Baseline Agent gpt-4-0613 $65.8$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $77.3$ $0.070$ gpt-3.5-turbo-16k-0613 $33.8$ $0.048$ MINT (Wang et al., 2024b): code subset, 136 instances MINT Baseline Agent gpt-4-0613 $59.6$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $50.0$ $0.087$ gpt-3.5-turbo-16k-0613 $5.2$ $0.030$ ProofWriter (Tafjord et al., 2021), 600 instances Few-shot Prompting + Chain-of-Thought (Pan et al., 2023) gpt4 $68.1$ - Logic-LM (Pan et al., 2023) gpt4 + symbolic solver $79.6$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $78.8$ - Entity Deduction Arena (Zhang et al., 2024a), 200 instances Zero-shot Prompting (Zhang et al., 2024a) gpt-4-0314 $40.0$ - gpt-3.5-turbo-0613 $27.0$ - OH CodeActAgent v1.5 gpt-4o-2024-05-13 $38.0$ - gpt-3.5-turbo-16k-0613 $24.0$ -
@@
AgentBench (Liu et al., 2023) evaluates agents’ reasoning and decision-making abilities in a multi-turn, open-ended generation setting. We selected the code-grounded operating system (OS) subset with 144 tasks. Agents from OpenHands interact directly with the task-specific OS using bash commands in a multi-turn manner, combining interaction and reasoning to automate task completion.
@@
MINT (Wang et al., 2024b) is a benchmark designed to evaluate agents’ ability to solve challenging tasks through multi-turn interactions using tools and natural language feedback simulated by GPT-4. We use coding and math subsets used in Yuan et al. (2024). We follow the original paper and allow the agent to interact with up to five iterations with two chances to propose solutions.
@@
ProofWriter (Tafjord et al., 2021) is a synthetic dataset created to assess deductive reasoning abilities of LLMs. Same as Logic-LM (Pan et al., 2023), we focus on the most challenging subset, which contains 600 instances requiring 5-hop reasoning. To minimize the impact of potential errors in semantic parsing, we use the logical forms provided by Logic-LM.
@@
Entity Deduction Arena (EDA) (Zhang et al., 2024a) evaluates agents’ ability to deduce unknown entities through strategic questioning, akin to the 20 Questions game. This benchmark tests the agent’s state tracking, strategic planning, and inductive reasoning capabilities over multi-turn conversations. We evaluate two datasets “Things” and “Celebrities”, each comprising 100 instances, and report the average success rate over these two datasets.